kyaunie:

12x-8-12? does that equal -180?

supie:

What is the value of x?

Imagine:

Why don't you subtract -8-12 first?

supie:

Yeah do that.

Imagine:

Putting equal terms together.

kyaunie:

okk

Imagine:

So, 12x-20=-180

Imagine:

12x-20=-180 +20 +20 12x=-160

Imagine:

\[\frac{ 12x }{ 12x }-\frac{ -160 }{ 12 }=13.3333333333\]

umm:

First step: simplifying. We start out with the equation below.\[12x-8-12\]So,\[=12x+-8+-12\]Second step: combining like terms\[=12x+−8+−12\]\[=(12x)+(−8+−12)\]\[=12x+−20\]Final answer:\[=12x+−20\]

Imagine:

Solve or Simplify?

Imagine:

@umm

umm:

Simplifying.